OpenStudy (anonymous):

How do you calculate x if: 0,97^x = 0,3

terenzreignz (terenzreignz):

Only logarithms... \[\huge a^x = b \rightarrow x = \log_a b\]

OpenStudy (anonymous):

hmm, how do I type that on a calculator?

OpenStudy (anonymous):

u have to use a scientific calculator..... or u may use a log table also...

OpenStudy (anonymous):

how? I can type log b, for instance, but how to I express the base a? I've got a ti-86

terenzreignz (terenzreignz):

Well, you have ln in that calculator?

terenzreignz (terenzreignz):

Or a log. It doesn't even matter, actually :D \[\huge \log_ab = \frac{\ln b}{\ln a}=\frac{\log b}{\log a}\]